# Deeper Insights into Deep Graph Convolutional Networks: Stability and Generalization

### v1

@@ A.2 Proof of Lemma 2
holds for any $k=1,2,\dots,K$. This completely proves (26). Furthermore, for the final output,
$|f(\mathbf{x}|\theta)-f(\mathbf{x}|\theta^{\prime})|=$ $|\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big{)}-\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big{)}|$
$\leq$ $\alpha_{\sigma}\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\big{(}\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big{)}\|_{F}$
$\leq$ $\alpha_{\sigma}\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}\|\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\|_{F}$
$\leq$ $\alpha_{\sigma}C_{g}\big{(}\|\mathbf{X}^{(K)}\|_{F}\|\triangle\mathbf{w}\|_{2}+\|\triangle\mathbf{X}^{(K)}\|_{F}\|\mathbf{w}^{\prime}\|_{2}\big{)}$
$\leq$ $B^{K}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*},$
which completes the proof of (27). ∎
Finally, for any $\mathbf{z}=(\mathbf{x},y)$ taken from $\mathcal{D}$, we denote by $\hat{y}=f(\mathbf{x}|\theta_{T})$ and $\hat{y}^{\prime}=f(\mathbf{x}|\theta^{\prime}_{T})$. Then, according to (27),
$\sup_{\mathcal{S},\mathbf{z}}\Big{|}\mathbb{E}_{\mathcal{A}}\big{[}\ell(\hat{y},y)\big{]}-\mathbb{E}_{\mathcal{A}}\big{[}\ell(\hat{y}^{\prime},y)\big{]}\Big{|}=$ $\sup_{\mathcal{S},z}\Big{|}\mathbb{E}_{\mathcal{A}}\big{[}\ell\big{(}f(\mathbf{x}|\theta_{T}),y\big{)}-\ell\big{(}f(\mathbf{x}|\theta^{\prime}_{T}),y\big{)}\big{]}\Big{|}$
$\leq$ $\alpha_{\ell}\sup_{\mathbf{x}}\mathbb{E}_{\mathcal{A}}\Big{[}\big{|}f(\mathbf{x}|\theta_{T})-f(\mathbf{x}|\theta^{\prime}_{T})\big{|}\Big{]}$
$\leq$ $\alpha_{\ell}B^{K}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\cdot\mathbb{E}_{\mathcal{A}}\big{[}\|\triangle\theta_{T}\|_{*}\big{]}.$
@@
A.3 Proof of Lemma 3 and Lemma 4
@@
$\big{\|}\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)$ $-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime})\big{\|}_{F}\leq\Big{(}\upsilon_{\sigma}B^{2K}\alpha^{2K}_{\sigma}C^{2K+2}_{g}C^{2}_{\mathbf{X}}+B^{K-1}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\Big{)}\|\triangle\theta\|_{*},$ (28)
@@
$\leq$ $\Big{(}\nu_{\sigma}B^{2K}\alpha^{2K}_{\sigma}C^{2K+2}_{g}C^{2}_{\mathbf{X}}+B^{K-1}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\Big{)}\|\triangle\theta\|_{*}+\rho_{k}\|\triangle\theta\|_{*},$ (29)
@@
$\rho_{k}:=\nu_{\sigma}(B\alpha_{\sigma}C_{g})^{K+k-1}C_{g}^{2}C^{2}_{\mathbf{X}}\Big{(}\sum_{j=0}^{K-k}(B\alpha_{\sigma}C_{g})^{j}\Big{)}.$ (30)
Proof. First, according to the proof of (26) and (27), the following holds true for $k=1,2,\dots,K+1$:
@@
$\leq$ $B^{k-1}\alpha^{k-1}_{\sigma}C^{k-1}_{g}C_{\mathbf{X}}\Big{(}\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)},$ (31)
where $\mathbf{W}^{(K+1)}=\mathbf{w}$. Furthermore,
$\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime})\|_{F}$
$=$ $\Big{\|}\nabla\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big{)}[g(\mathbf{L})\mathbf{X}^{(K)}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}-\nabla\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big{)}[g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\Big{\|}_{F}$
@@
$+\Big{\|}\nabla\sigma\big{(}\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big{)}[g(\mathbf{L})\triangle\mathbf{X}^{(K)}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\Big{\|}_{F}$
$\leq$ $\upsilon_{\sigma}|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}-\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}|\cdot\|\mathbf{X}^{(K)}\|_{F}\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}+\alpha_{\sigma}\|\triangle\mathbf{X}^{(K)}\|_{F}\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}$
$\leq$ $\upsilon_{\sigma}C_{g}\|\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\|_{F}\cdot\|\mathbf{X}^{(K)}\|_{F}\cdot C_{g}+\alpha_{\sigma}C_{g}\|\triangle\mathbf{X}^{(K)}\|_{F}.$
Combining (23), (26) and (31),
@@
which completes the proof of (28). Next, we turn to prove (29). First, for $k=1,2,\dots,K$,
@@
By (23), (24) and (26), we have
$\leq$ $B^{K-1}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\Big{(}\sum_{j=1}^{k-1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)}+B^{k-1}\alpha_{\sigma}^{k-1}C_{g}^{k}C_{\mathbf{X}}\cdot\gamma_{k},$ (32)
where $\gamma_{k}:=\big{\|}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}-\frac{\partial f(\mathbf{x}|\theta^{\prime})}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)^{\prime}}\big{\|}_{F}$. Now, we need to bound $\gamma_{k}$.
@@
where $h_{k}:=\big{\|}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\big{(}\mathbf{R}^{(k)}-\mathbf{R}^{(k)^{\prime}}\big{)}\big{\|}_{F}$. By (31),
@@
$\leq$ $\nu_{\sigma}B^{k-1}\alpha^{k-1}_{\sigma}C^{k}_{g}C_{\mathbf{X}}\Big{(}\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)}.$
Combining (24), we have
$h_{k}$ $=\big{\|}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\big{(}\mathbf{R}^{(k)}-\mathbf{R}^{(k)^{\prime}}\big{)}\big{\|}_{F}\leq\big{\|}\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\big{\|}_{F}\|\mathbf{R}^{(k)}-\mathbf{R}^{(k)^{\prime}}\|_{F}$
$\leq\nu_{\sigma}B^{K}\alpha^{K}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\Big{(}\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)}.$
It is easy to see that
$h_{k}\leq h_{k+1}\leq\dots\leq h_{K}\leq\nu_{\sigma}B^{K}\alpha^{K}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*}.$
@@
$\gamma_{k}\leq h_{K}+\alpha^{2}_{\sigma}C_{g}(B\alpha_{\sigma}C_{g})^{K-k}\|\triangle\mathbf{W}^{(k+1)}\|_{2}+B\alpha_{\sigma}C_{g}\cdot\gamma_{k+1}.$
@@
$\leq$ $h_{K}+\alpha^{2}_{\sigma}C_{g}\|\triangle\mathbf{w}\|_{2}+\nu_{\sigma}B^{K+1}\alpha^{K+1}_{\sigma}C^{K+2}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*}.$
@@
$\gamma_{k}\leq$ $h_{K}\Big{(}\sum_{j=0}^{K-k}(B\alpha_{\sigma}C_{g})^{j}\Big{)}+\alpha^{2}_{\sigma}C_{g}(B\alpha_{\sigma}C_{g})^{K-k}\Big{(}\sum_{j=k+1}^{K+1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)}$
@@
$\leq$ $h_{K}\Big{(}\sum_{j=0}^{K-k}(B\alpha_{\sigma}C_{g})^{j}\Big{)}+\alpha^{2}_{\sigma}C_{g}(B\alpha_{\sigma}C_{g})^{K-k}\Big{(}\sum_{j=k+1}^{K+1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)}$
$+\nu_{\sigma}B^{2K+1-k}\alpha^{2K+1-k}_{\sigma}C^{2K+2-k}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*},$ (33)
where $\triangle\mathbf{W}^{(K+1)}=\triangle\mathbf{w}$. Finally, substituting (33) into (32),
$\leq$ $B^{K-1}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\Big{(}\sum_{j=1}^{k-1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big{)}$
$+B^{k-1}\alpha_{\sigma}^{k-1}C_{g}^{k}C_{\mathbf{X}}\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}-\frac{\partial f(\mathbf{x}|\theta^{\prime})}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)^{\prime}}\|_{F}$
@@
which completes the proof of (29).

### v2

@@ Acknowledgment
holds for any $k=1,2,\dots,K$. This completely proves (A.8).
Furthermore, for the final output, using the Lipschitz property of $\sigma(\cdot)$, we have
$|f(\mathbf{x}|\theta)-f(\mathbf{x}|\theta^{\prime})|=|\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)-\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)|\leq\alpha_{\sigma}|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\big(\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)|.$
Note that
$|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\big(\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)|\leq\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}\cdot\|\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\|_{F}\leq C_{g}\big(\|\mathbf{X}^{(K)}\|_{F}\|\triangle\mathbf{w}\|_{2}+\|\triangle\mathbf{X}^{(K)}\|_{F}\|\mathbf{w}^{\prime}\|_{2}\big).$
Combining (A.5) and (A.8), we further have
$|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\big(\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)|\leq B^{K}\alpha_{\sigma}^{K}C_{g}^{K+1}C_{\mathbf{X}}(\|\triangle\mathbf{w}\|_{2}+\sum_{j=1}^{K}\|\triangle\mathbf{W}^{(j)}\|_{2})=B^{K}\alpha_{\sigma}^{K}C_{g}^{K+1}C_{\mathbf{X}}\|\triangle\theta\|_{*}.$
Thus,
$|f(\mathbf{x}|\theta)-f(\mathbf{x}|\theta^{\prime})|\leq\alpha_{\sigma}|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\big(\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)|\leq B^{K}\alpha_{\sigma}^{K+}C_{g}^{K+1}C_{\mathbf{X}}\|\triangle\theta\|_{*},$
which completes the proof of (A.9).
Proof of Lemma 2: Now, we are ready to prove Lemma 2 based on Lemma 6. For any $\mathbf{z}=(\mathbf{x},y)$ taken from $\mathcal{D}$, we denote by $\hat{y}=f(\mathbf{x}|\theta_{T})$ and $\hat{y}^{\prime}=f(\mathbf{x}|\theta^{\prime}_{T})$. Firstly, using the Lipschitz property of loss function $\ell(\cdot,\cdot)$, we have
$\sup_{\mathcal{S},\mathbf{z}}\Big|\mathbb{E}_{\mathcal{A}}\big[\ell(\hat{y},y)\big]-\mathbb{E}_{\mathcal{A}}\big[\ell(\hat{y}^{\prime},y)\big]\Big|=\sup_{\mathcal{S},z}\Big|\mathbb{E}_{\mathcal{A}}\big[\ell\big(f(\mathbf{x}|\theta_{T}),y\big)-\ell\big(f(\mathbf{x}|\theta^{\prime}_{T}),y\big)\big]\Big|\leq\alpha_{\ell}\sup_{\mathbf{x}}\mathbb{E}_{\mathcal{A}}\Big[\big|f(\mathbf{x}|\theta_{T})-f(\mathbf{x}|\theta^{\prime}_{T})\big|\Big]$
Then, according to (A.9),
$\sup_{\mathcal{S},\mathbf{z}}\Big|\mathbb{E}_{\mathcal{A}}\big[\ell(\hat{y},y)\big]-\mathbb{E}_{\mathcal{A}}\big[\ell(\hat{y}^{\prime},y)\big]\Big|\leq\alpha_{\ell}B^{K}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\cdot\mathbb{E}_{\mathcal{A}}\big[\|\triangle\theta_{T}\|_{*}\big].$
@@
APPENDIX C: Proof of Lemma 3 and Lemma 4
@@
$\big\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime})\big\|_{F}\leq\Big(\upsilon_{\sigma}B^{2K}\alpha^{2K}_{\sigma}C^{2K+2}_{g}C^{2}_{\mathbf{X}}+B^{K-1}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\Big)\|\triangle\theta\|_{*},$ (A.10)
@@
$\big\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime})\big\|_{F}\leq\Big(\nu_{\sigma}B^{2K}\alpha^{2K}_{\sigma}C^{2K+2}_{g}C^{2}_{\mathbf{X}}+B^{K-1}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\Big)\|\triangle\theta\|_{*}+\rho_{k}\|\triangle\theta\|_{*},$ (A.11)
@@
$\rho_{k}:=\nu_{\sigma}(B\alpha_{\sigma}C_{g})^{K+k-1}C_{g}^{2}C^{2}_{\mathbf{X}}\Big(\sum_{j=0}^{K-k}(B\alpha_{\sigma}C_{g})^{j}\Big).$ (A.12)
Proof. First, according to the proof of (A.8) and (A.9), the following holds true for $k=1,2,\dots,K+1$:
@@
$\leq$ $B^{k-1}\alpha^{k-1}_{\sigma}C^{k-1}_{g}C_{\mathbf{X}}\Big(\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big),$ (A.13)
where $\mathbf{W}^{(K+1)}=\mathbf{w}$.
We now prove (A.10). First, applying $\mathbf{A}_{1}\mathbf{A}_{2}-\mathbf{A}^{\prime}_{1}\mathbf{A}^{\prime}_{2}=(\mathbf{A}_{1}-\mathbf{A}_{1}^{\prime})\mathbf{A}_{2}+\mathbf{A}^{\prime}_{1}(\mathbf{A}_{2}-\mathbf{A}_{2}^{\prime})$, we have
$\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime})\|_{F}=$ $\Big\|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)[g(\mathbf{L})\mathbf{X}^{(K)}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}-\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)[g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\Big\|_{F}$
@@
$+\Big\|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)[g(\mathbf{L})\triangle\mathbf{X}^{(K)}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\Big\|_{F}.$
Using the $\nu_{\sigma}$-smooth property of $\sigma(\cdot)$ and applying $\|\mathbf{A}_{1}\mathbf{A}_{2}\|_{F}\leq\|\mathbf{A}_{1}\|_{2}\|\mathbf{A}_{2}\|_{F}$, we have
$\Big\|\Big(\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)-\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)\Big)[g(\mathbf{L})\mathbf{X}^{(K)}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\Big\|_{F}$
$\leq$ $\Big|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big)-\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)\Big|\cdot\Big\|[g(\mathbf{L})\mathbf{X}^{(K)}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\Big\|_{F}$
$\leq$ $\upsilon_{\sigma}|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}-\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}|\cdot\|\mathbf{X}^{(K)}\|_{F}\|\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\|_{2}$
$\leq$ $\upsilon_{\sigma}C_{g}\|\mathbf{X}^{(K)}\mathbf{w}-\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\|_{F}\cdot\|\mathbf{X}^{(K)}\|_{F}\cdot C_{g},$
and since $|\nabla\sigma(\cdot)|\leq\alpha_{\sigma}$, $\Big\|\nabla\sigma\big(\boldsymbol{\delta}_{\mathbf{x}}^{\top}g(\mathbf{L})\mathbf{X}^{(K)^{\prime}}\mathbf{w}^{\prime}\big)[g(\mathbf{L})\triangle\mathbf{X}^{(K)}]^{\top}\boldsymbol{\delta}_{\mathbf{x}}\Big\|_{F}\leq\alpha_{\sigma}C_{g}\|\triangle\mathbf{X}^{(K)}\|_{F}$. Then, combining (A.5), (A.8) and (A.13), we have
@@
which completes the proof of (A.10).
Next, we turn to prove (A.11). First, for $k=1,2,\dots,K$,
@@
Let
$\gamma_{k}:=\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)}-\frac{\partial f(\mathbf{x}|\theta^{\prime})}{\partial\mathbf{X}^{(k)}}\odot\mathbf{R}^{(k)^{\prime}}\big\|_{F}.$ (A.14)
Then, combining (A.5), (A.6) and (A.8), we have
$\big\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime})\big\|_{F}\leq B^{K-1}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\Big(\sum_{j=1}^{k-1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big)+B^{k-1}\alpha_{\sigma}^{k-1}C_{g}^{k}C_{\mathbf{X}}\cdot\gamma_{k},$ (A.15)
Next, we need to bound $\gamma_{k}$.
@@
where $h_{k}:=\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\big(\mathbf{R}^{(k)}-\mathbf{R}^{(k)^{\prime}}\big)\big\|_{F}$. By (A.13),
@@
$\leq$ $\nu_{\sigma}B^{k-1}\alpha^{k-1}_{\sigma}C^{k}_{g}C_{\mathbf{X}}\Big(\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big).$ (A.16)
Combining (A.6), we have
$h_{k}$ $=\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\odot\big(\mathbf{R}^{(k)}-\mathbf{R}^{(k)^{\prime}}\big)\big\|_{F}\leq\big\|\frac{\partial f(\mathbf{x}|\theta)}{\partial\mathbf{X}^{(k)}}\big\|_{F}\cdot\|\mathbf{R}^{(k)}-\mathbf{R}^{(k)^{\prime}}\|_{F}$
$\leq\nu_{\sigma}B^{K}\alpha^{K}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\Big(\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big).$ (A.17)
Let $h_{\max}=\nu_{\sigma}B^{K}\alpha^{K}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*}$. Then, it is easy to see that
$h_{k}\leq h_{\max}~\textrm{holds for all}~k=1,2,\dots,K.$ (A.18)
@@
$\gamma_{k}\leq h_{\max}+\alpha^{2}_{\sigma}C_{g}(B\alpha_{\sigma}C_{g})^{K-k}\|\triangle\mathbf{W}^{(k+1)}\|_{2}+B\alpha_{\sigma}C_{g}\cdot\gamma_{k+1}.$
@@
$\leq$ $h_{\max}+\alpha^{2}_{\sigma}C_{g}\|\triangle\mathbf{w}\|_{2}+\nu_{\sigma}B^{K+1}\alpha^{K+1}_{\sigma}C^{K+2}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*}.$
@@
$\gamma_{k}\leq$ $h_{\max}\Big(\sum_{j=0}^{K-k}(B\alpha_{\sigma}C_{g})^{j}\Big)+\alpha^{2}_{\sigma}C_{g}(B\alpha_{\sigma}C_{g})^{K-k}\Big(\sum_{j=k+1}^{K+1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big)$
@@
$\leq$ $h_{\max}\Big(\sum_{j=0}^{K-k}(B\alpha_{\sigma}C_{g})^{j}\Big)+\alpha^{2}_{\sigma}C_{g}(B\alpha_{\sigma}C_{g})^{K-k}\Big(\sum_{j=k+1}^{K+1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big)$
$+\nu_{\sigma}B^{2K+1-k}\alpha^{2K+1-k}_{\sigma}C^{2K+2-k}_{g}C_{\mathbf{X}}\|\triangle\theta\|_{*},$ (A.19)
where $\triangle\mathbf{W}^{(K+1)}=\triangle\mathbf{w}$. Finally, substituting (A.19) into (A.15),
$\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime})\|_{F}\leq$ $B^{K-1}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\Big(\sum_{j=1}^{k-1}\|\triangle\mathbf{W}^{(j)}\|_{2}\Big)+B^{k-1}\alpha_{\sigma}^{k-1}C_{g}^{k}C_{\mathbf{X}}\cdot\gamma_{k}$
$\leq$ $\Big(\nu_{\sigma}B^{2K}\alpha^{2K}_{\sigma}C^{2K+2}_{g}C^{2}_{\mathbf{X}}+B^{K-1}\alpha_{\sigma}^{K+1}C_{g}^{K+1}C_{\mathbf{X}}\Big)\|\triangle\theta\|_{*}$
@@
which completes the proof of (A.11).